1. You jog at a speed of 6 miles per hour for 20 minutes, and you walk at a speed of 4 miles per hour for t minutes. If your combined jogging and walking distance is 3 miles, for how long did you walk?
Answer by stanbon(75887)

You jog at a speed of 6 miles per hour for 20 minutes, and you walk at a speed of 4 miles per hour for t minutes. If your combined jogging and walking distance is 3 miles, for how long did you walk?
---
Jogging DATA:
rate = 6 mph ; time = 1/3 hr ; distance = tr = 2 miles
---------------------------
Walking DATA:
rate = 4 mph : time = t/60 hr ; distance = tr = t/15 miles
-------------------------
Equation:
miles + miles = 3 miles
2 miles + t/15 = 3
t/15 = 1
t = 15 minutes
t = (1/4) hour
=======================
Cheers,
Stan H.